1. Renz is approaching a stoplight moving with a velocity of +30.0 m/s. The light turns yellow, and Renz applies the brakes and skids to a stop. If Renz acceleration is -8.00 m/s2, then determine the displacement of the car during the skidding process.

2. A car slows down its motion from 10 m/s to 6 m/s in 2 seconds under constant acceleration. What is its acceleration?

3. A car starts its motion from rest and uniformly accelerates at a rate of 4 m/s2 for 2 seconds in a straight line. How far did the car travel during those 2 seconds?

4. An object, without change in direction, travels a distance of 50 m with an initial speed of 5 m/s in 4 s. Find the object's velocity at the end of the given time interval? 

5. Ben Simmons is waiting at a stoplight. When it finally turns green, Ben accelerated from rest at a rate of 6.00 m/sfor a time of 4.10 seconds. Determine the displacement of Ben's car during this time period.



 

 

 

 

 

 

 

 

 

 



1
Expert's answer
2021-12-06T09:42:40-0500

1)

"v^2=v_0^2+2ad,""0=v_0^2+2ad,""d=-\dfrac{v_0^2}{2a}=-\dfrac{(30\ \dfrac{m}{s})^2}{2\times(-8\ \dfrac{m}{s^2})}=56.25\ m."

2)

"v=v_0+at,""a=\dfrac{v-v_0}{t}=\dfrac{6\ \dfrac{m}{s}-10\ \dfrac{m}{s}}{2\ s}=-2\ \dfrac{m}{s}."

3)

"d=v_0t+\dfrac{1}{2}at^2,""d=0+\dfrac{1}{2}\times4\ \dfrac{m}{s^2}\times(2\ s)^2=8\ m."

4)

"d=\dfrac{1}{2}(v_0+v)t,""v=\dfrac{2d}{t}-v_0=\dfrac{2\times50\ m}{4\ s}-5\ \dfrac{m}{s}=20\ \dfrac{m}{s}."

5)

"d=v_0t+\dfrac{1}{2}at^2,""d=0+\dfrac{1}{2}\times6\ \dfrac{m}{s^2}\times(4.1\ s)^2=50.43\ m."